Solve for x and y :-
 
x/a^2 + y/b^2 = 2
 
x/a + y/b = a + b 
 
Thanks. 

Grade:10

5 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
Given Equations are
x/a^2 + y/b^2 = 2..........(1)
x/a + y/b = a + b...........(2)
Simplifying equations (1) and (2) by taking lcm we get
b2x+a2y=2a2b2..........(4)
bx+ay=a2b+ab2........(3)
Multiplying equation 3 with b we get
b2x+aby=a2b2+ab3..........(5)
b2x+a2y=2a2b2............(6)
subtracting 5 from 6 we get
y=b2and substituting y=b2inx/a + y/b = a + b we get x=a2
so x=a2and y=b2

Siddharth Gurla
9 years ago
I did not understand the Subtracting (5) from (6) part.
SHAIK AASIF AHAMED
9 years ago

Hello student,
Please find the answer to your question below
Given Equations are
x/a^2 + y/b^2 = 2..........(1)
x/a + y/b = a + b...........(2)
Simplifying equations (1) and (2) by taking lcm we get
b2x+a2y=2a2b2..........(4)
bx+ay=a2b+ab2........(3)
Multiplying equation 3 with b we get
b2x+aby=a2b2+ab3..........(5)
b2x+a2y=2a2b2............(6)
subtracting 5 from 6 we get
(ab-a2)y=ab3-a2b2
a(b-a)y=ab2(b-a)
After cancelling a(b-a) on both sides we get
y=b2
substituting y=b2in x/a + y/b = a + b we get x=a2
I think you will understand it now
Siddharth Gurla
9 years ago
Dear Sir , 
 
Just wanted to let you know that I have understood the solution, 
SHAIK AASIF AHAMED
9 years ago
Hello student,
We understand your needs and value your appreciation. We always work to make the study easy foryou. Your feedbacks are valuable for us so please keep posting your feedbacks and suggestions tohelp us in further improving our quality.